# When Hyperspectral Image Classification Meets Diffusion Models: An Unsupervised Feature Learning Framework

### v1

@@ III Method
Our proposed Diff-HSI aims to learn effective spectral-spatial representations in an unsupervised manner via denoising diffusion probabilistic model (DDPM) for HSI classification. The framework is shown in Fig. 2. Our method consists of two key steps. The first step is to pretrain the DDPM with unlabeled patches to learn complex spectral-spatial distributions and dependencies of HSI data. After pretraining, we freeze the parameters of the pretrained DDPM. The second step of our Diff-HSI is a supervised learning process aiming to fully extract and dynamically fuse the timestep-wise features from the pretrained denoising U-Net for classification. In order to capture the characteristics of different objects in HSI data, we first extract abundant spectral-spatial features from different timestep $t$ to construct the timestep-wise feature bank. Then, we design a dynamic feature fusion module that fuses the features in the feature bank dynamically and generates multi-timestep representations with discriminative information for classification. Finally, the representations are fed into an ensemble of lightweight classifiers to predict the label of the center pixel. In this section, we introduce our proposed Diff-HSI in detail, including the unsupervised diffusion-based pretraining step as well as the supervised finetuning step with multi-timestep representation generation.
III-A Diffusion-based Unsupervised Spectral-Spatial Feature Learning
In order to learn complex spectral-spatial relations and label-agnostic information of HSI data, we pretrain a diffusion model in an unsupervised manner in the first step of our Diff-HSI. We first give a brief review of DDPM and introduce the basic principles to make our method easier to understand. Then, we explain our unsupervised feature learning procedure in detail through diffusion-based pretraining with HSI data.
@@ III-A1 A Brief Review of DDPM
Denoising diffusion probabilistic models are a class of likelihood-based models that reconstruct the distribution of training data via an encoder-decoder denoising model. The denoising model is trained to remove noise from the training data destructed by Gaussian noises step-by-step. These models consist of a forward noising process and a reverse denoising process. In the forward process, Gaussian noise is added to the original training data $x_{0}\sim q\left(x_{0}\right)$ step by step over $T$ time steps, which follows the Markovian process:
@@
Figure 3: 2D t-SNE visualization of features from different timesteps on the Indian Pines dataset. (a) original spectral feature. (b) feature at $t=100$. (c) feature at $t=300$. (d) feature at $t=500$. (e) feature at $t=700$. (f) feature at $t=900$.
@@ III-A2 Unsupervised Hyperspectral Diffusion Pretraining
Using the training skills and optimization objectives mentioned above, our DDPM is trained by unlabeled hyperspectral data. Suppose that $S$ is the set of all samples in a dataset, $D$ is a denoising diffusion probabilistic model, and $E$ is an ensemble of lightweight classifiers. The target of the unsupervised pretraining step is to learn label-agnostic spectral-spatial features on unlabelled data sampled from $S$ so that the features can be leveraged in the finetuning step to train ensemble lightweight classifiers using a small amount of labeled data sampled from $S_{labeled}\in S$ for classification.
@@
In order to learn effective spectral-spatial features in an unsupervised manner, we first pretrain a denoising diffusion probabilistic model with unlabeled patches randomly cropped from the HSI dataset. Before training, the data is pre-processed by principal components analysis (PCA) and patch cropping operation. Then, patches are randomly sampled from compressed HSI for DDPM pretraining. In detail, given an unlabeled patch $x_{0}\in\mathcal{R}^{H\times H}$, we gradually add Gaussian noise to the HSI patch according to the variance schedule $\{\beta_{t}\}_{t=0}^{T}$ in the diffusion process where $T$ is the total number of the timestep. Then, in the reverse process, a denoising U-Net is trained to predict the noise added on $x_{t-1}$ taking noisy patch $x_{t}$ and timestep $t$ as inputs. And $x_{t-1}$ is calculated by subtracting the predicted noise from $x_{t}$. In other words, the model tries to capture some useful information of $x_{t}$ to reconstruct $x_{t-1}$ that is closer to $x_{0}$. In each step of the training process, the timestep $t$ is randomly sampled from 0 to $T$.
In the pretraining procedure with unlabeled hyperspectral data, the DDPM is able to automatically learn the label-agnostic prior of HSI data that reveals the complex dependencies and relations in the spectral and spatial dimensions. At different timesteps, different input noisy patches and their optimization goals allow the DDPM to attend to varied information of the data, which is demonstrated to be critical to HSI classification.
III-B Multi-Timestep Representation for Supervised Finetuning
After unsupervised pretraining, in order to explore the rich information learned by the pretrained DDPM, the second step of our method is to extract effective features from the pretrained DDPM and construct multi-timestep representations for HSI classification. We first extract abundant features and construct a timestep-wise feature bank. Then, a dynamic feature fusion module is designed to automatically fuse the features into multi-timestep representations. Finally, ensemble lightweight classifiers are trained for classification.
III-B1 Timestep-Wise Feature Bank
In the finetuning step, the parameters of the pretrained DDPM are frozen. The denoising U-Net is able to capture meaningful information from the input data in the reverse process. Leveraging the rich and diverse information captured in the reverse steps is crucial for accurate classification. Therefore, timestep-wise spectral-spatial features are extracted from the intermediate hierarchies of DDPM at different timesteps to construct effective representations that contain diverse information of input HSI. Given an input patch $x_{0}\in\mathcal{R}^{H\times H}$ randomly cropped from HSI and a timestep $t$, Gaussian noise is gradually added to $x_{0}$ according to the variance schedule through diffusion process $q(x_{t}|x_{0})$. The noisy patch $x_{t}$ can be directly obtained by equation 2. Then, the noisy patch $x_{t}$ is fed into the pretrained denoising U-Net to obtain hierarchal features from the U-Net decoder, which implies rich spectral information as well as multi-scale spatial information, as shown in Fig. 2. The features from different layers of the decoder are jointly upsampled to $H\times H$ and then concatenated to get the feature $f_{t}\in\mathcal{R}^{H\times H\times d}$ at timestep $t$. The above process is defined as a function $\mathcal{F}$ and formulated as follows:
$f_{t}=\mathcal{F}(D,x_{t},t)$ (6)
where $D$ is the pretrained denoising U-Net.
@@ III-B1 Timestep-Wise Feature Bank
For each feature $f_{t_{i}}\in\mathcal{R}^{H\times H\times d}$, we reserve only the vector corresponding to the center pixel indexed as $(\frac{H}{2},\frac{H}{2})$ to obtain the pixel-wise feature $h_{t_{i}}\in\mathcal{R}^{1\times 1\times d}$, which largely reduce computational cost with fewer parameters. Following the above process, we draw $m$ sets of features to construct the timestep-wise feature bank:
=$\mathcal{B}=\{h_{t_{i}}|i\in\{1,...,m\},h_{t_{i}}=center(f_{t_{i}})\}$ (7)
where $t_{1},...,t_{m}$ are sampled from $[0,T]$ at equal intervals. Features extracted at different timestep $t$ tend to capture different information of HSI data since the reverse process is a step-by-step denoising process that tries to predict further information conditioned on the given destructed data at every different timestep.
To further demonstrate this, we visualize the features in 2-dimensional space using t-distributed stochastic neighbor embedding (TSNE) to show the diversity of features at different timesteps $t$. The visualization of the original spectral feature space is shown in Fig. 3(a), and the features in the timestep-wise feature bank at timestep 100, 300, 500, 700, and 900 are visualized in Fig. 3(b)-(f) respectively. As observed, the features in the timestep-wise feature bank show great diversity. Furthermore, with the increase of timestep $t$, the clusters of the features tend to be more loosely structured, and features of the same class tend to be formed into a few large clusters rather than dispersed into multiple smaller clusters. Features at smaller $t$ tend to capture stochastic variation, and features at larger $t$ tend to focus on high-level information, both of which are key to HSI classification with different categories of land covers. Thus the abundant features in the feature bank contain diverse and multi-level information of the input HSI data, exhibiting diverse properties and concerns.
III-B2 Dynamic Feature Fusion Module
For each input HSI patch $x_{0}$, a timestep-wise feature bank is obtained with abundant features containing varied characteristics of the patch. However, the features contain a wealth of information, but not all of it is essential. Using all the features for classification will bring high computational costs and redundant information, and manually selecting timestep-wise features from the feature bank is difficult to characterize spectral-spatial features effectively and robustly. Meanwhile, HSI data have different spectral characteristics for different HSI datasets, resulting in different spectral representation laws of timestep-wise features. Also, data on the same dataset is very different due to its position. Therefore, the dynamic feature fusion module is crucial to be proposed, and is capable of adaptively fusing features of different timesteps from the DDPM to learn effective multi-timestep representations with key information.
To obtain dynamic multi-timestep representations, we design the dynamic feature fusion module to predict the soft dynamic weight of multi-timestep representations. The architecture of the dynamic feature fusion module is illustrated in Fig. 2. Specifically, Let $L=[h_{t_{1}}...h_{t_{m}}]$ denotes the concatenated features from the timestep-wise feature bank as the input of the dynamic weight network. The dynamic weight network consists of two linear layers and outputs the $n$ sets of dynamic weights $w$ with a SoftMax function:
$w$ $=\mathcal{M}(L,W)=SoftMax(a(L,W))$ (8)
$=SoftMax(W_{2}\delta(W_{1}L))$
where $M$ is the dynamic weight network, $W$ is the linear weight of $M$, $\delta$ is ReLU function, $w\in\mathcal{R}^{m\times 1\times 1\times n}$ and $L\in\mathcal{R}^{m\times 1\times 1\times n}$. Inspired by the design of the multi-head attention mechanism in [42], we generate multiple sets of dynamic weights, which allows the representations to focus on different information from different positions and spectral bands. Each dynamic weight determines the proportion of each feature in the corresponding representation. Then, the dynamic weights are used to weight the pixel-wise features by matrix multiplication to obtain $n$ multi-timestep representations:
$r_{i}=w^{T}_{i}L$ (9)
where $r_{i}$ is the i-th representation, $i\in{1,...,n}$. The multi-step representations will be concatenated into the final representation to be classified. Thus, the most appropriate and relevant representation is automatically learned from the feature bank without human intervention, saving time cost and reducing the impact of redundant information. $n$ will be ablated to find the most suitable value balancing the performance and computational cost.
III-B3 Lightweight Ensemble Classifiers
After getting the dynamic representation, a lightweight network is needed to predict the classification label. Inspired by [43], we train an ensemble of lightweight linear classifiers that takes the dynamic pixel representations as inputs and predicts the classification label of each pixel. Specifically, each classifier is trained independently, consisting of two hidden layers with ReLU activation and batch normalization. When testing a sample, the final predicted label is obtained by majority voting of the ensemble of pixel classifiers, as illustrated in Fig. 2. This method brings more stability of prediction with a very small cost since the parameters of each classifier are very limited.
IV Experiments and results
In this section, we first describe the three used well-known HSI datasets, including the Indian Pines dataset, Pavia University dataset and Houston2018 dataset. The experimental setting is then introduced including evaluation metrics, a brief introduction of compared state-of-art methods, and implementation details. Then, we conduct quantitative experiments and ablation analysis to evaluate our proposed method.
IV-A Datasets Description
IV-A1 Indian Pines
IV-A2 Pavia University
The University of Pavia dataset was collected in 2003 by the reflective optics system imaging spectrometer (ROSIS-3) sensor over a part of the city of Pavia, Italy. The dataset consists of $610\times 340$ pixels with a spatial resolution of 1.3 m and 115 spectral bands in the wavelength range of 430 to 860 nm. 103 out of 115 bands are used for classification after removing 12 noisy bands. The image contains a large number of background pixels and only 42776 labeled pixels divided into 9 classes, including asphalt, meadows, gravel, and so on. We use 5% of the labeled samples for training and the rest for testing. The class name and the number of training and testing samples are listed in Table II.
IV-A3 Houston2018
The Houston2018 dataset, identified as 2018 IEEE GRSS DFC dataset, was gathered in 2018 by the National Center for Airborne Laser Mapping (NCALM) over the University of Houston campus and its neighboring urban area, including HSI, multispectral LiDAR, and very high resolution RGB images. The HSI dataset consists of $601\times 2384$ pixels with a spatial resolution of 1 m and 48 spectral bands in the wavelength range of 380 to 1050 nm. It contains 504856 labeled pixels and 20 classes of interest. We use 5% of the labeled samples for training and the rest for testing. The class name and the number of training and testing samples are listed in Table III.

### v2

@@ III Method
Our proposed MTMSD is a novel diffusion-based feature learning framework and aims to explore multi-timestep multi-stage diffusion features effectively and efficiently for modeling spectral-spatial relations comprehensively. The framework is shown in Fig. 2. In the following section, we introduce the proposed MTMSD in detail.
III-A Diffusion Feature Learning on unlabeled HSI
@@ III-A1 A Brief Review of DDPM
DDPMs are a class of likelihood-based models that reconstruct the distribution of training data via an encoder-decoder denoising model. The denoising model is trained to remove noise from the training data destructed by Gaussian noises step-by-step. These models consist of a forward noising process and a reverse denoising process. In the forward process, Gaussian noise is added to the original training data $x_{0}\sim q\left(x_{0}\right)$ step by step over $T$ time steps, which follows the Markovian process:
@@ III-A2 Unsupervised Hyperspectral Diffusion Pretraining
Using the training skills and optimization objectives mentioned above, the DDPM for diffusion feature learning is trained with unlabeled hyperspectral data. Before training, the HSI data is pre-processed by principal components analysis (PCA) and random patch cropping operation. Then, given an unlabeled patch $x_{0}\in\mathcal{R}^{H\times H\times D}$, where $H$ is the patch size, and $D$ is the number of PCA components, we gradually add Gaussian noise to the unlabeled HSI patch according to the cosine variance schedule $\{\beta_{t}\}_{t=0}^{T}$ in the diffusion process, where $T$ is the total number of the timestep. Then, in the reverse process, a denoising U-Net is trained to predict the noise added on $x_{t-1}$ taking noisy patch $x_{t}$ and timestep $t$ as inputs. And the sample $x_{0}$ can be obtained from the noise patch $x_{t}$ by the iterative denoising steps according to Eq. 3. In each step of the training process, the timestep $t$ is randomly sampled from 0 to $T$. The U-Net denoising model $\epsilon_{\theta}(x_{t},t)$ is optimized by minimizing Eq. 5. The parameters in the pretraining process, such as patch size, number of PCA components, and total pretraining steps, will be discussed in Sec IV. F.
Figure 3: Purification index generation of our proposed class & timestep-oriented feature purification module.
III-B Multi-Timestep Multi-Stage Diffusion Feature Extraction
After diffusion feature learning on unlabeled HSI, there exist abundant multi-timestep multi-stage diffusion features that contain both contextual semantics and textual information hidden in the denoising U-Net with different timestep $t$. To model the complex spectral-spatial relations in HSIs, we extract multi-timestep multi-stage features from all stages of the fixed pretrained denoising U-Net decoder with all timesteps, and construct the timestep-wise center and global feature bank by center extraction and average pooling.
Specifically, given a labeled patch $x_{0}^{{}^{\prime}}\in\mathcal{R}^{H\times H\times D}$ preprocessed by PCA to $D$ channels, $x_{0}^{{}^{\prime}}$ is corrupted by adding Gaussian noise according to Eq. 2 and obtain $\{x^{{}^{\prime}}_{t_{i}}\}_{i=1}^{m}$ at a set of timesteps $\{t_{i}\}_{i=1}^{m}$ that are sampled from $[0,T]$ at equal intervals. Then, the noisy patches $\{x^{{}^{\prime}}_{t_{i}}\}_{i=1}^{m}$ are fed into the pretrained denoising U-Net to extract multi-timestep multi-stage diffusion features from all stages of the U-Net decoder. The different layer features are jointly upsampled to $H\times H$ and then concatenated to get the multi-stage feature $f_{t_{i}}\in\mathcal{R}^{H\times H\times d}$ at timestep $t_{i}$.
@@ III-B Multi-Timestep Multi-Stage Diffusion Feature Extraction
For each multi-stage feature $f_{t_{i}}$, we only reserve the center feature $c_{t_{i}}\in\mathcal{R}^{1\times 1\times d}$ located as $(\frac{H}{2},\frac{H}{2})$ corresponding to the center pixel, and obtain the global feature $g_{t_{i}}\in\mathcal{R}^{1\times 1\times d}$ through global average pooling, which largely reduce computational cost with fewer memories. Following the above process, we construct the timestep-wise center feature bank $\mathcal{B}_{c}$ and the timestep-wise global feature bank $\mathcal{B}_{g}$:
$\mathcal{B}_{c}=\{c_{t_{i}}|i\in\{1,...,m\},c_{t_{i}}=center(f_{t_{i}})\}$ (6)
$\mathcal{B}_{g}=\{g_{t_{i}}|i\in\{1,...,m\},g_{t_{i}}=avgpool(f_{t_{i}})\}$ (7)
III-C Class & Timestep-oriented Multi-Stage Feature Purification
Multi-stage features, extracted from the denoising U-Net, embody abundant reconstruction information from the pretraining process of the diffusion model. Despite this richness, the information is not entirely aligned with HSI classification requirements, containing redundant features irrelevant to the task. Furthermore, these features exhibit a degree of repetition among themselves. Therefore, the class & timestep-oriented multi-stage purification is proposed to explore multi-stage features by selecting significant channels, aiming to remove the redundant information and reduce the computational cost.
Before the multi-stage feature purification, we first generate the purification index using the prior of dataset-wise multi-timestep multi-stage feature bank, depicted in Fig. 3. Specifically, for a $C$-category HSI classification dataset, $S^{j}$ is the training samples of category $j$, $j\in\{1,...,C\}$. The feature banks of samples in $S^{j}$ are averaged to get the category representative features of category $j$. Representing all training samples by the category representative features can significantly reduce computational overhead. Gathering all the representative features of each category, a representative feature matrix $M\in\mathcal{R}^{m\times C\times d}$ can be obtained,
$M_{i,j}=\frac{1}{|S^{j}|}\sum_{x\in S^{j}}c_{t_{i}}(x)$ (8)
where $c_{t_{i}}(x)$ is the feature at $t_{i}$ in the center feature bank of sample $x$, $|S^{j}|$ is the size of $S^{j}$.
In order to purify the discriminative and effective channels for classification, a class-oriented significance score is proposed to evaluate the significance of each channel from inter-class relations. It aims to filter out channels that are highly homogenized and thus have little impact on classification by minimizing the inter-class similarity and maximizing the inter-class variance. For $k^{th}$ index, the class-oriented significance score $\tau_{class}^{k}$ is formed as:
$\tau_{class}^{k}=-\alpha U_{class}^{k}+(1-\alpha)V_{class}^{k}$ (9)
where $\alpha\in(0,1)$, $U_{class}^{k}$ denotes the inter-class similarity at index $k$ obtained by summing the average cosine similarities across classes at all the timesteps, and $V_{class}^{k}$ denotes the inter-class variance at index $k$ which is the sum of the variances across classes at all the timesteps, formed as follows.
$U_{class}^{k}=\frac{1}{m}\frac{1}{c^{2}}\sum_{i=1}^{m}\sum_{p=1}^{c}\sum_{\begin{subarray}{c}q=1\\ q\neq p\end{subarray}}^{c}m_{i,p,k}\cdot m_{i,q,k}$ (10)
$V_{class}^{k}=\frac{1}{m}\frac{1}{c}\sum_{i=1}^{m}\sum_{p=1}^{c}(M_{i,p,k}-\frac{1}{c}\sum_{q=1}^{c}M_{i,q,k})^{2}$ (11)
Similarly, to preserve the diversity of features while reducing repetitive information at the timestep dimension different from the class dimension, the timestep-oriented significance score $\tau_{t}^{k}$ at index $k$ is designed to be calculated as:
$\tau_{t}^{k}=-\beta U_{t}^{k}+(1-\beta)V_{t}^{k}$ (12)
where $\beta\in(0,1)$, $U_{t}^{k}$ denotes the inter-timestep similarity at index $k$ obtained by summing the average cosine similarities across timesteps at all the classes, and $V_{t}^{k}$ is the inter-timestep variance at index $k$ which is the sum of the variances across classes at all the timesteps, defined as follows.
$U_{t}^{k}=\frac{1}{c}\frac{1}{m^{2}}\sum_{i=1}^{c}\sum_{p=1}^{m}\sum_{\begin{subarray}{c}q=1\\ q\neq p\end{subarray}}^{m}M_{p,i,k}\cdot M_{q,i,k}$ (13)
$V_{t}^{k}=\frac{1}{C}\frac{1}{m}\sum_{i=1}^{c}\sum_{p=1}^{m}(m_{p,i,k}-\frac{1}{m}\sum_{q=1}^{m}m_{q,i,k})^{2}$ (14)
Finally, the class & timestep-oriented significance score of each index $k\in\{1,...,d\}$ is obtained as the final measurement.
$\tau^{k}=\tau_{class}^{k}+\tau_{t}^{k}$ (15)
We rank the indexes by the class & timestep-oriented significance score $\tau^{k}$ and generate the purification index by reserving the indexes of the top $K$ highest $\tau^{k}$, which indicates the most inter-class and inter-timestep divergence and discrimination. Then, multi-stage features in the timestep-wise center feature bank $\mathcal{B}_{c}$ multiply the purification index to obtain the purified timestep-wise center feature bank $\mathcal{B}_{c}^{{}^{\prime}}$.
Figure 4: The structure of global feature-guided selective timestep network in our proposed selective timestep feature fusion.
III-D Selective Timestep Feature Fusion
For each labeled HSI patch $x_{0}^{{}^{\prime}}$, multi-timestep features contain textual information and contextual semantics from shallow to deep timesteps. Meanwhile, HSI data have different spectral characteristics for different HSI datasets, resulting in different spectral representation laws of timestep-wise features. Compared to the manual selection of single timestep feature in previous works, we explore the multi-timestep features and perform selective timestep feature fusion to adaptively select different timestep features integrating textual information and contextual semantics and learn comprehensive multi-timestep multi-stage selective representations. The whole process is illustrated in Fig. 2.
Specifically, for the purified timestep-wise center feature bank $\mathcal{B}_{c}^{{}^{\prime}}=\{c_{t_{i}}^{{}^{\prime}}|i\in\{1,...,m\}\}$, our goal is to adaptively select the timestep of features in it for modeling spectral-spatial relations better. We first fuse all purified center features in $\mathcal{B}_{c}^{{}^{\prime}}$ to obtain the all-timestep feature $c_{t}\in\mathcal{R}^{1\times 1\times K}$ via an element-wise summation:
$c_{t}=\sum_{i=1}^{m}c_{t_{i}}^{{}^{\prime}}$ (16)
Then, the all-timestep feature $c_{t}$ is fed into a simple multi-layer perception (MLP) to obtain a compact feature $z\in\mathcal{R}^{1\times 1\times K_{r}}$ with fewer channels for better efficiency:
$z=\mathcal{F}_{mlp}(c_{t})=W_{1}(\delta(BN(W_{2}c_{t})))$ (17)
where $\delta$ is the ReLU function, $BN$ is the batch normalization, $W_{1}\in\mathcal{R}^{K\times K_{r}}$, $W_{2}\in\mathcal{R}^{K_{r}\times K_{r}}$, $K_{r}=K/2$.
However, adaptive timestep selection solely based on timestep-wise center features leads to imprecise choices due to the absence of spatial information from neighboring pixels around the central pixel. Hence, we propose a global feature-guided selective timestep network, designed to incorporate global features for guiding multi-timestep feature selection, thereby enriching spatial information in modeling spatial distributions. And the network is depicted in Fig. 4. Specifically, given a timestep-wise global feature bank $\mathcal{B}_{g}=\{g_{t_{i}}|i\in\{1,...,m\}\}$, for each purified center feature $c_{t_{i}}^{{}^{\prime}}$, the corresponding global feature $g_{t_{i}}$ is concatenated with $z$ with the transformation $\mathcal{F}_{i}$ to obtain $h_{i}$. Then the linear projection $\mathcal{W}_{i}$ is applied to the $h_{i}$ respectively, $i\in\{1,...,m\}$. To adaptively select different timesteps of center features, a softmax operator on the channel-wise digits is used to obtain the selective weights $\{w_{i}\}_{i=1}^{m}$ guided by the compact feature $z$ and the corresponding global information:
$w^{{}^{\prime}}_{i}=\mathcal{W}(h_{i})=\mathcal{W}([\mathcal{F}_{i}(z),g_{t_{i}}])$ (18)
$w^{c}_{i}=\frac{e^{w^{{}^{\prime}c}_{i}}}{\sum_{j=1}^{m}e^{w^{{}^{\prime}c}_{j}}}$ (19)
where $\{\mathcal{F}_{i}\}_{i=1}^{m}$ and $\{\mathcal{W}_{i}\}_{i=1}^{m}$ are linear projections to align the channel dimension to $K$ channels, and $w^{c}_{i}$ is the $c$-th element of the selective weight $w_{i}$, $c\in\{1,...,K\}$. Finally, the multi-timestep multi-stage selective representation $r_{s}$ is selected and fused through the selective weights $\{w_{i}\}_{i=1}^{m}$ on each purified center feature $c_{t_{i}}^{{}^{\prime}}$:
$r_{s}=\sum_{i=1}^{m}w_{i}c_{t_{i}}^{{}^{\prime}}$ (20)